A 100 kg bumper-car at a fairground (carA) is rolling along a level surface in the positive x-direction at a speed of 1.00 m/s. It collides with a second 120kg car (car B) which is moving in the same direction at a speed of 0.10 m/s. If the cars stick together after the collision, what is their common velocity?
Answer #1

Let the common velocity be V m/s

By the conservation of momentum we have

m1u1 + m2u2 = (m1 + m2)V

100*1 + 120*0.1 = (100 + 120)V

V = 0.51 m/s

so their common velocity is 0.51 m/s